# Searching Latent Program Spaces

### v2

@@ 4.2 Search Methods for Latent Optimization
This means we search for the latent that would most likely make the decoder generate the right outputs given the corresponding inputs. By finding a latent that can explain all the input-output pairs, the latent solution to the optimization problem is more likely to generalize to a new input-output pair. We describe here two instantiations of the search process, namely a sampling and a gradient ascent algorithm, both acting in the latent space of programs. We leave for future work the exploration of other search methods like evolutionary strategies (Hansen and Ostermeier, 2001; Chalumeau et al., 2023) that could better trade-off exploration and exploitation of the latent space.
@@
Since the decoder is a differentiable neural network, its log-likelihood $\log p_{\theta}(y|x,z)$ is also differentiable with respect to $z$ and one can use first-order methods like gradient-ascent to efficiently search through the latent space for a solution to the latent optimization problem (equation 4, see Figure 18 for a visualization of a 2D latent space trained on grid patterns). This visualization highlights that only a small portion of the latent space can explain all the input-output pairs, corresponding to high decoding likelihood. Notably, poor initialization can lead the search to converge to different local minima, highlighting the importance of amortized inference from the encoder. We initialize $z_{0}^{\prime}$ as the mean of all the pair latents sampled from the encoder, and refine it iteratively as follows:
@@ 4.3 Training
To train the LPN system end-to-end, we assume we have a dataset of tasks, where a task is defined as $n$ input-output pairs $(x_{i},y_{i})$ generated by the same program. To simulate the test conditions of predicting a new input from a given specification, we design the training procedure to reconstruct each of the outputs $y_{i}$ from their inputs $x_{i}$ and all the $n-1$ other pairs ${(x_{j},y_{j})}_{j\neq i}$. We emphasize that we do not use the specific pair $(x_{i},y_{i})$ to reconstruct $y_{i}$, which would lead to the encoder directly compressing the output $y_{i}$ as a shortcut without learning program-related abstractions.
When reconstructing output $y_{i}$, we first sample latents $z_{j}$ from the encoder $q_{\phi}(z|x_{j},y_{j})$ for all $j\neq i$. We then aggregate them by computing their mean $\frac{1}{n-1}\sum_{j\neq i}z_{j}$, then we perform latent optimization using e.g., gradient ascent to obtain $z_{i}^{\prime}$. Finally, we compute the negative log-likelihood of the right output $y_{i}$ using its corresponding input $x_{i}$ and the refined latent $z_{i}^{\prime}$. In practice, we compute the cross-entropy loss of the decoder logits $p_{\theta}(\hat{y_{i}}|x_{i},z_{i}^{\prime})$ and the labels $y_{i}$, which is derived from maximizing the likelihood of a categorical distribution. The full training pipeline is detailed in Appendix C. Specifically, we compute the reconstruction loss $\mathcal{L_{\text{rec}}}$ and the KL loss $\mathcal{L_{\text{KL}}}$ between the approximate posterior and the prior:
@@
The dependence of the reconstruction loss $\mathcal{L_{\text{rec}}}(\phi,\theta)$ in $\phi$ arises from using the reparameterization trick (Kingma, 2013) when sampling each latent $z_{j}$. Indeed, we first sample a normal random vector $\epsilon_{j}\sim\mathcal{N}(0,I)$, then we infer the mean $\mu_{j}$ and diagonal covariance $\Sigma_{j}$ using the encoder and recompose the latent $z_{j}=\mu_{j}+\epsilon_{j}\cdot\Sigma_{j}$. Then, $z_{i}^{\prime}$ is used by the decoder to reconstruct the output. Note that we can decide whether to let the decoder gradient flow through the latent update. Indeed, it is more computationally efficient to stop the gradient through the update, by changing line 9 of algorithm 2 with $z_{i}^{\prime}=z_{i}^{\prime}+\alpha\cdot\overline{g_{i}^{\prime}}$, where $g_{i}^{\prime}=\nabla_{z}\sum_{j\neq i}\log p_{\theta}(y_{j}|x_{j},z)|_{z=z_{i}^{\prime}}$, with $\overline{g_{i}^{\prime}}$ notating a stop-gradient on $g_{i}^{\prime}$.
We denote $\beta$ as the weighting factor that balances the reconstruction and KL terms (Burgess et al., 2018), which gives the combined training objective: $\mathcal{L}_{\text{total}}(\phi,\theta)=\mathcal{L}_{\text{rec}}(\phi,\theta)+\beta\mathcal{L}_{\text{KL}}(\phi)$. This training procedure offers some freedom in the latent optimization, i.e., how to compute $z_{i}^{\prime}$ from $z_{i}$. Training with gradient ascent latent optimization (as detailed in Algorithm 2) incurs a compute overhead due to the cost of the latent gradient computation through the decoder. Although we may use a high compute budget at test-time, we propose to use a small number of gradient ascent steps during training, ranging from 0 to 5 steps.
@@ 5 Experiments
Table 1: Ablation of LPN training and inference methods on the Pattern task. Rows/columns represent different training/inference methods, differing only in the latent optimization. "Grad [N]" stands for N gradient ascent steps, "Sample [N]" for N samples from the encoder distribution without leveraging any gradients, and "Grad 1 ⋆⋆" means that the decoder parameter gradient flows through the latent optimization update. Training was performed for 20k steps with 3 seeds, aggregating performance as mean (and standard deviation in brackets) over the 3 runs. Bold values indicate the best training method for each inference regime. See expanded table in Section B.1
@@ 5.1 Setup
We consider the ARC-AGI 2024 challenge (Chollet et al., 2024) as the testing domain for our method. This program synthesis benchmark encompasses diverse tasks designed to test adaptation and out-of-distribution generalization, rather than memorization. We implement both the LPN encoder and decoder as small transformers (Vaswani et al., 2017), see Appendix G for full architecture details. We fully open-source the codebase used to run all experiments in the paper in the supplementary material. We introduce the simpler Pattern task to investigate LPN’s dynamics before large-scale ARC-AGI training, see Appendix A. It generates 10x10 black input grids with a blue pixel indicating where a 4x4 program-specific pattern should be pasted. This task is sufficient to demonstrate weaknesses in deep learning models that do not leverage test-time computation (see Section 5.4).
@@
We compare LPN to an in-context learning method Kolev et al. (2020); Li et al. (2024a), which encodes each of the input-output pairs and then concatenates these embeddings to condition output prediction, notably never producing any intermediate program embeddings. We discuss the motivation for this design in Appendix H. Then, we also compare to test-time fine-tuning (Devlin et al., 2017a; Akyürek et al., 2024), which, given a test-time specification, performs parameter-based gradient updates on the in-context model. See Appendix H for details on baseline implementations and Appendix E for hyperparameters.
@@ 5.2 Pattern Task
We compare a variety of LPN training and inference methods in Table 1, to better understand the dynamics of LPN. We aim to first answer whether LPN can self-improve at test-time by searching its latent space, and then how the training inference strategy affects performance. For each training method, we train a small 1M parameter model for 20k steps with a batch size of 128 and evaluate it with different inference modes. We find that inference using no test-time adaptation performs poorly across all instances of LPN (and in-context baselines, see Section B.2). This shows that amortizing inference is difficult for models of this size and training time.
All variations of LPN training show strong scaling in performance as the number of gradient steps at test time is scaled. This demonstrates that LPN is capable of test-time adaptation to improve pre-trained amortization performance. We visualise the decoded outputs at many points in the latent space of LPN on a simple pattern task in Section B.4 showing structure in the latent program representations, and Section B.4 shows the likelihood of the specification is smooth across the space, enabling gradient based test time search.
When comparing different LPN training inference strategies, 1 gradient ascent step of latent optimization shows higher returns than training with mean pooling, with the difference particularly pronounced when scaling the inference budget. With 100 steps of gradient ascent at inference time, LPN training with grad 0 gets an accuracy of 67.5%, whereas training with one gradient step (Grad 1) reaches 99.5%. This demonstrates the benefits of training the latent space with the awareness that gradient ascent will be performed at test-time, an important inductive bias for the LPN architecture. We also observe that gradient ascent vastly outperforms sampling-based search, validating that a search without any gradient signal is highly inefficient. Lastly we perform training with and without the stop gradient on the gradient update itself. We found that using the stop gradient actually leads to higher test time performance while also being more computationally efficient.
@@ 5.3 Starting test-time search from the encoder
Figure 2: Ablation on the role of the encoder and latent optimization. Latents are initialized from the mean of the encoder latents (except for orange). Grad $N$ stands for doing latent optimization with $N$ gradient steps. Both the encoder initialization and the latent optimization matter for LPN.
In addition, we ablate the impact of initializing latent optimization with the encoder versus using the prior, in order to assess the benefits of amortizing inference. In Figure 2, we specifically compare no latent search (blue), to doing latent optimization in two ways: sampling from the prior $z\sim p(z)$ (orange), and sampling from the encoder $z\sim q_{\phi}(z|x,y)$ (green). Initializing latent search with the encoder is critical for performance across all training methods. This validates the intuition that LPN can perform fast system 1-like reasoning using the encoder and then narrow down the search space during latent optimization, simulating system 2 reasoning.
@@ 5.4 Adapting Out-Of-Distribution
Table 2: Out-of-distribution (OOD) performance on the Pattern task. Performance is averaged over 3 training runs with different seeds, with standard deviation in parentheses.
A significant challenge to deep learning methods is generalizing to out-of-distribution (OOD) tasks. We study the OOD behavior of LPN, in-context learning, and parameter fine-tuning on the Pattern task in Table 2. In-context learning has no test time adaptation performance as it has no inherent mechanism to scale compute beyond a single forward pass. We include full experiments on a range of varying OOD settings in Section B.2. When trained with 1 step of gradient ascent, LPN recovers high accuracy (88%) using 100 steps of gradient ascent at inference time. TTT and in-context learning are incapable of recovering the right OOD pattern at inference time. We perform a similar experiment in Section B.7 on a synthetic integer sequence task with less spatial structure, with the same conclusions.
@@ 5.5 Scaling Specification
Figure 3: Scaling specification sizes on the OOD pattern task.
For any learning method that can handle varying amounts of data, an important question is how the method scales to handling more data than it was trained on. In Figure 3 we evaluate in-context learning, TTT, and LPN with and without gradient search in the OOD setting of the pattern task, with varying specification sizes. LPN smoothly generalizes to different specification sizes, with only a small drop in performance with less data, and keeps improving with more data. In-context learning overfits to the specification size it is trained on, and TTT only performs well once given a sufficiently large specification and high inference compute. We ablate this in Section B.5, repeating the experiment for varying training specification sizes.
@@ 5.6 ARC-AGI 2024
We train on the re-arc dataset (Hodel, 2024), designed to be in distribution relative to the ARC training set (which we don’t train on) Li et al. (2024b). The evaluation set is significantly OOD, representing a challenging generalization experiment. We train a 178M-parameter LPN with a 256-dim latent space for 100k steps for 2 days on a TPU v4-32, see Appendix E. Leveraging insights from pattern task experiments, we use gradient ascent in LPN during training. To remain efficient, we train LPN in Grad 0 mode for 95k steps and then fine-tune in Grad 1 for a further 5k steps, which we find gives a marginal boost in performance at lower levels of test time compute Section B.8.
@@ Results
Table 3 shows top-2 accuracy on ARC-AGI (standard to ensure resolvability of all problems). Even after significant amortization of inference, we observe large performance gains from scaling gradient search at test time. LPN significantly outperforms TTT in-distribution. We also see strong evidence that the LPN has learned a structured latent space for representing programs, demonstrated by a t-SNE plot of the embeddings of different programs in the training set, see Appendix F. Out of distribution, LPN doubles its performance by using latent space search at test-time (scaling FLOPs from 2e11 to 2e15). Using from 2E+11 to 2E+13 FLOPs, LPN outperforms TTT, until we scale test-time compute to much higher levels, where TTT achieves higher performance. This finding is likely due to the lack of diversity in programs seen by LPN during training, relative to experiments on other datasets, limiting the ability to have a smooth expressive latent space. Relative to the pattern task, TTT is less susceptible to overfitting on ARC-AGI, possibly due to the use of larger model sizes. Both methods outperform two ARC specific LLM approaches that leverage far larger pre-trained models, including CodeT5 (Butt et al., 2024) (220M parameters) and text-davinci (Mirchandani et al., 2023) (175B parameters). This work represents an improvement over such methods due to significantly lower inference and training cost, and greater generalization from limited training distribution. The current highest performing model on ARC-AGI is o3-preview(low) (LLM) which achieves 75.7% at a cost of \$200 per task. Both the training and inference costs are far beyond this work, there is an opportunity for future work to investigate scaling up this work to such a scale.
@@
Table 3: Results on ARC-AGI for varying values of FLOPs. In-Distribution refers to puzzles in the ARC train set not trained on but are in-distribution relative to re-ARC. Out-of-distribution refers to puzzles from the eval set. LPN and TTT stand for Latent Program Network (ours), and Test-Time Training. CodeIt and Mirchandani are baselines from prior work.

### v3

@@ 4.2 Search Methods for Latent Optimization
This means we search for the latent that would most likely make the decoder generate the right outputs given the corresponding inputs. By finding a latent that can explain all the input-output pairs, the latent solution to the optimization problem is more likely to generalize to a new input-output pair. We describe here two instantiations of the search process, namely a sampling and a gradient ascent algorithm, both acting in the latent space of programs. We leave for future work the exploration of other search methods like evolutionary strategies (cmaes; chalumeau2023combinatorial) that could better trade-off exploration and exploitation of the latent space.
@@
Since the decoder is a differentiable neural network, its log-likelihood $\log p_{\theta}(y|x,z)$ is also differentiable with respect to $z$ and one can use first-order methods like gradient-ascent to efficiently search through the latent space for a solution to the latent optimization problem (equation 4, see Figure˜20 for a visualization of a 2D latent space trained on grid patterns). This visualization highlights that only a small portion of the latent space can explain all the input-output pairs, corresponding to high decoding likelihood. Notably, poor initialization can lead the search to converge to different local minima, highlighting the importance of amortized inference from the encoder. We initialize $z_{0}^{\prime}$ as the mean of all the pair latents sampled from the encoder, and refine it iteratively as follows:
@@ 4.3 Training
To train the LPN system end-to-end, we assume we have a dataset of tasks, where a task is defined as $n$ input-output pairs $(x_{i},y_{i})$ generated by the same program. To simulate the test conditions of predicting a new input from a given specification, we design the training procedure to reconstruct each of the outputs $y_{i}$ from their inputs $x_{i}$ and all the n1n-1 other pairs ${(x_{j},y_{j})}_{j\neq i}$. We emphasize that we do not use the specific pair $(x_{i},y_{i})$ to reconstruct $y_{i}$, which would lead to the encoder directly compressing the output $y_{i}$ as a shortcut without learning program-related abstractions.
When reconstructing output $y_{i}$, we first sample latents $z_{j}$ from the encoder $q_{\phi}(z|x_{j},y_{j})$ for all $j\neq i$. We then aggregate them by computing their mean $\frac{1}{n-1}\sum_{j\neq i}z_{j}$, then we perform latent optimization using e.g., gradient ascent to obtain $z_{i}^{\prime}$. Finally, we compute the negative log-likelihood of the right output $y_{i}$ using its corresponding input $x_{i}$ and the refined latent $z_{i}^{\prime}$. In practice, we compute the cross-entropy loss of the decoder logits $p_{\theta}(\hat{y_{i}}|x_{i},z_{i}^{\prime})$ and the labels $y_{i}$, which is derived from maximizing the likelihood of a categorical distribution. The full training pipeline is detailed in Appendix˜C. Specifically, we compute the reconstruction loss $\mathcal{L_{\text{rec}}}$ and the KL loss $\mathcal{L_{\text{KL}}}$ between the approximate posterior and the prior:
@@
The dependence of the reconstruction loss $\mathcal{L_{\text{rec}}}(\phi,\theta)$ in $\phi$ arises from using the reparameterization trick (kingma2013auto) when sampling each latent $z_{j}$. Indeed, we first sample a normal random vector $\epsilon_{j}\sim\mathcal{N}(0,I)$, then we infer the mean $\mu_{j}$ and diagonal covariance $\Sigma_{j}$ using the encoder and recompose the latent $z_{j}=\mu_{j}+\epsilon_{j}\cdot\Sigma_{j}$. Then, $z_{i}^{\prime}$ is used by the decoder to reconstruct the output. Note that we can decide whether to let the decoder gradient flow through the latent update. Indeed, it is more computationally efficient to stop the gradient through the update, by changing line 9 of algorithm 2 with $z_{i}^{\prime}=z_{i}^{\prime}+\alpha\cdot\overline{g_{i}^{\prime}}$, where $g_{i}^{\prime}=\nabla_{z}\sum_{j\neq i}\log p_{\theta}(y_{j}|x_{j},z)|_{z=z_{i}^{\prime}}$, with $\overline{g_{i}^{\prime}}$ notating a stop-gradient on $g_{i}^{\prime}$.
We denote $\beta$ as the weighting factor that balances the reconstruction and KL terms (burgess2018understanding), which gives the combined training objective: $\mathcal{L}_{\text{total}}(\phi,\theta)=\mathcal{L}_{\text{rec}}(\phi,\theta)+\beta\mathcal{L}_{\text{KL}}(\phi)$. This training procedure offers some freedom in the latent optimization, i.e., how to compute $z_{i}^{\prime}$ from $z_{i}$. Training with gradient ascent latent optimization (as detailed in Algorithm˜2) incurs a compute overhead due to the cost of the latent gradient computation through the decoder. Although we may use a high compute budget at test-time, we propose to use a small number of gradient ascent steps during training, ranging from 0 to 5 steps.
@@ 5 Experiments
Table 1: Ablation of LPN training and inference methods on the Pattern task, reporting exact match accuracy (%). Rows/columns represent different training/inference methods, differing only in the latent optimization. Grad [N] stands for NN gradient ascent steps, Sample [N] for NN samples from the encoder distribution without leveraging any gradients, and Grad 1 ⋆⋆ means that the decoder parameter gradient flows through the latent optimization update. Training was performed for 20k steps with 3 seeds, aggregating performance as mean (and standard deviation in brackets) over the 3 runs. Bold values indicate the best training method for each inference regime. See expanded table in Section˜B.1.
@@ 5.1 Setup
We consider the ARC-AGI 2024 challenge (arc-prize-2024) as the testing domain for our method. This program synthesis benchmark encompasses diverse tasks designed to test adaptation and out-of-distribution generalization, rather than memorization. We implement both the LPN encoder and decoder as small transformers (vaswani2017attention), see Appendix˜G for full architecture details. We introduce the simpler Pattern task to investigate LPN’s dynamics before large-scale ARC-AGI training, see Appendix˜A. It generates 10x10 black input grids with a blue pixel indicating where a 4x4 program-specific pattern should be pasted. This task is sufficient to demonstrate weaknesses in deep learning models that do not leverage test-time computation (see Section˜5.5).
@@
We compare LPN to an in-context learning method kolev2020neural; li2024combining, which encodes each of the input-output pairs and then concatenates these embeddings to condition output prediction, notably never producing any intermediate program embeddings. We discuss the motivation for this design in Appendix˜H. Then, we also compare to test-time fine-tuning (devlin2017neural; akyurek2024surprising), which, given a test-time specification, performs parameter-based gradient updates on the in-context model. See Appendix˜H for details on baseline implementations and Appendix˜E for hyperparameters.
@@ 5.2 Pattern Task
We compare a variety of LPN training and inference methods in Table˜1, to better understand the dynamics of LPN. We aim to first answer whether LPN can self-improve at test-time by searching its latent space, and then how the training inference strategy affects performance. For each training method, we train a small 1M-parameter model for 20k steps with a batch size of 128 and evaluate it with different inference modes. We find that inference using no test-time adaptation performs poorly across all instances of LPN (and in-context baselines, see Section˜B.2). This shows that amortizing inference is difficult for models of this size and training time.
All variations of LPN training show strong scaling in performance as the number of gradient steps at test time is scaled. This demonstrates that LPN is capable of test-time adaptation to improve pre-trained amortization performance. We visualise the decoded outputs at many points in the latent space of LPN on a simple pattern task in Section˜B.4 showing structure in the latent program representations, and Section˜B.4 shows the likelihood of the specification is smooth across the space, enabling gradient based test time search.
When comparing different LPN training inference strategies, 1 gradient ascent step of latent optimization shows higher returns than training with mean pooling, with the difference particularly pronounced when scaling the inference budget. With 100 steps of gradient ascent at inference time, LPN training with Grad 0 gets an accuracy of $67.5\%$, whereas training with one gradient step (Grad 1) reaches $99.5\%$. This demonstrates the benefits of training the latent space with the awareness that gradient ascent will be performed at test-time, an important inductive bias for the LPN architecture. We also observe that gradient ascent vastly outperforms sampling-based search, validating that a search without any gradient signal is highly inefficient. Lastly, we perform training with and without the stop gradient on the gradient update itself. We found that using the stop gradient actually leads to higher test time performance while also being more computationally efficient.
5.3 String Manipulation Task
To investigate the generalizability of our results beyond environments with significant spatial structure, we perform an ablation on a synthetic sequence task and replicate the analysis previously conducted on the pattern dataset.
Inference
Training Grad 0 Grad 10 Grad 100
In-Context 77.7 (1.2) - -
TTT 78.1 (1.1) 72.8 (0.8) 11.8 (1.9)
LPN Grad 0 75.6 (1.7) 83.3 (0.8) 81.3 (0.5)
LPN Grad 1 71.3 (0.4) 86.1 (1.5) 81.8 (1.8)
LPN Grad 1 ⋆⋆ 67.9 (1.3) 85.0 (0.4) 81.0 (1.3)
Table 2: Exact match accuracy (%) on the test set for the string manipulation task. The comparison includes an In-Context baseline, Test-Time Training (TTT) (lr=1e4lr=1e-4), and three LPN variants (with lr=1e1lr=1e-1): Grad 0, Grad 1, and Grad 1 ⋆⋆ (decoder gradient flows through latent optimization). Performance is averaged over 3 runs, with standard deviations in parentheses.
This synthetic dataset features a vast program space, with over 100 million unique programs, each defined by composing 3 to 5 parameterized rules that transform sequences of numbers (ranging from 0 to 4), see Section˜B.7 for further details. Our experiments demonstrate that Test-Time Training (TTT) exhibits overfitting, as evidenced by its performance degradation when gradient fine-tuning is performed. In contrast, LPN maintains robust performance without overfitting. Furthermore, incorporating a single gradient step during training enhances LPN performance marginally at higher inference gradients.
@@ 5.4 Starting test-time search from the encoder
Figure 2: Ablation on the role of the encoder and latent optimization. Latents are initialized from the mean of the encoder latents (except for orange). Grad NN stands for doing latent optimization with $N$ gradient steps. Both the encoder initialization and the latent optimization matter for LPN.
We experiment on the Pattern task again, this time to assess the benefits of amortizing inference. We ablate the impact of initializing latent optimization with the encoder versus using the prior.
We specifically compare no latent search (blue) to doing latent optimization in two ways: sampling from the prior $z\sim p(z)$ (orange) and sampling from the encoder $z\sim q_{\phi}(z|x,y)$ (green).
Initializing latent search with the encoder is critical for performance across all training methods. This validates the intuition that LPN can perform fast system 1-like reasoning using the encoder and then narrow down the search space during latent optimization, simulating system 2 reasoning.
@@ 5.5 Adapting Out-Of-Distribution
Table 3: Out-of-distribution (OOD) performance on the Pattern task, measured as exact match accuracy (%). Performance is averaged over 3 training runs with different seeds, with standard deviation in parentheses.
A significant challenge to deep learning methods is generalizing to out-of-distribution (OOD) tasks. We study the OOD behavior of LPN, in-context learning, and parameter fine-tuning on the Pattern task in Table˜3. In-context learning has no test time adaptation performance as it has no inherent mechanism to scale compute beyond a single forward pass. We include full experiments on a range of varying OOD settings in Section˜B.2.
When trained with 1 step of gradient ascent, LPN recovers high accuracy (88%) using 100 steps of gradient ascent at inference time. TTT and in-context learning are incapable of recovering the right OOD pattern at inference time. We perform a similar experiment in Section˜B.7 on a synthetic integer sequence task with less spatial structure, with the same conclusions.
@@ 5.6 Scaling Specification
Figure 3: Exact match accuracy (%) on the out-of-distribution (OOD) pattern task as the specification size is scaled from 1 input output pair to 19 pairs, for different inference methods.
For any learning method that can handle varying amounts of data, an important question is how the method scales to handling more data than it was trained on. In Figure˜3 we evaluate in-context learning, TTT, and LPN with and without gradient search in the OOD setting of the pattern task, with varying specification sizes. LPN smoothly generalizes to different specification sizes, with only a small drop in performance with less data, and keeps improving with more data. In-context learning overfits to the specification size it is trained on, and TTT only performs well once given a sufficiently large specification and high inference compute. We ablate this in Section˜B.5, repeating the experiment for varying training specification sizes.
@@ 5.7 ARC-AGI 2024
We train on the re-arc dataset (hodel2024addressingabstractionreasoningcorpus), designed to be in distribution relative to the ARC training set (which we don’t train on) li2024tackling. The evaluation set is significantly OOD, representing a challenging generalization experiment. We train a 178M-parameter LPN with a 256-dim latent space for 100k steps for 2 days on a TPU v4-32, see Appendix˜E. Leveraging insights from pattern task experiments, we use gradient ascent in LPN during training. To remain efficient, we train LPN in Grad 0 mode for 95k steps and then fine-tune in Grad 1 for a further 5k steps, which we find gives a marginal boost in performance at lower levels of test time compute Section˜B.8.
@@ Results
Table˜4 shows top-2 accuracy on ARC-AGI (standard to ensure resolvability of all problems). Even after significant amortization of inference, we observe large performance gains from scaling gradient search at test time. LPN significantly outperforms TTT in-distribution. We also see strong evidence that the LPN has learned a structured latent space for representing programs, demonstrated by a t-SNE plot of the embeddings of different programs in the training set, see Appendix˜F. Out of distribution, LPN doubles its performance by using latent space search at test-time (scaling FLOPs from 2e11 to 2e15). Using from 2E+11 to 2E+13 FLOPs, LPN outperforms TTT, until we scale test-time compute to much higher levels, where TTT achieves higher performance. This finding is likely due to the lack of diversity in programs seen by LPN during training, relative to experiments on other datasets, limiting the ability to have a smooth expressive latent space. See Section˜B.11 for an ablation on the performance contribution to LPN from encoder initialization and gradient search.
Relative to the pattern task, TTT is less susceptible to overfitting on ARC-AGI, possibly due to the use of larger model sizes. Both methods outperform two ARC-specific LLM approaches that leverage far larger pre-trained models, including CodeT5 (butt2024codeit) (220M parameters) and text-davinci (mirchandani2023large) (175B parameters). This work represents an improvement over such methods due to significantly lower inference and training cost, and greater generalization from a limited training distribution. At the time of writing, the highest performing model on ARC-AGI is o3-preview(low) (LLM) which achieves 75.7% at a cost of \$200 per task. Both the training and inference costs are far beyond this work, there is an opportunity for future work to investigate scaling up this work to such a scale. To further contextualise the results compared to other methods applied to ARC-AGI we outline in Appendix˜I a complete table of leading and diverse approaches to the benchmark highlighting the training assumptions and compute for all methods where the information is public.
In section Section˜B.9 we investigate whether the problems solved, on the ARC-AGI evaluation test set, by TTT and LPN are different. We find that indeed different subsets of the test set are solved by each method and we visualize 3 examples of solutions only solved by LPN and TTT respectively.
Lastly, in Section˜B.10, we explore LPN’s capacity for compositional generalization at test time. We find that the latent search process can compose programs seen during training into novel combinations, which avoids the overfitting seen in models without search, that tend to execute only a single program. However, this capability is not robust; our analysis identifies both successful examples where the search enables composition and clear failure cases where it does not.
@@
Table 4: Results on ARC-AGI, measured by exact match accuracy (%), for varying values of FLOPs. In-Distribution refers to puzzles in the ARC train set , not included in the training set, but are in-distribution relative to re-ARC. Out-of-distribution refers to puzzles from the eval set. LPN and TTT stand for Latent Program Network (ours), and Test-Time Training. CodeIt and Mirchandani are baselines from prior work.